### v1

@@ 5 Conclusion
This supplementary material provides additional details about the proposed method NAVER. In particular, it provides details about the ablation studies for Answerer in section 6, the ablation studies for VLM in section 7, the analysis of self-correction mechanisms in section 8, the performance analysis for query length in section 9, and the prompts used for the LLMs and VLMs in section 10.
6 Ablation Studies for Answerer
We evaluate the impact of the Answering state in NAVER by assessing its role when explicit logic reasoning is either enabled or disabled. The results, presented in Table 10, demonstrate the significant contribution of the Answerer state to the system’s performance. When explicit logic reasoning is disabled (✗), Answerer validates each entity to determine if it satisfies the query. Incorporating Answerer improves the IoU scores from 67.2 to 80.8 on RefCOCO and from 70.3 to 76.9 on RefCOCO+. When explicit logic reasoning is enabled (✓), Answerer is used to validate the reasoned target candidates. This further increases performance, improving IoU scores from 70.8 to 83.3 on RefCOCO and from 73.1 to 78.4 on RefCOCO+. Answerer addresses potential errors occurring from Perception or Logic Generation by ensuring the final output aligns with the query. These findings highlight the importance of Answerer in refining predictions and enhancing robustness. By validating the outputs, Answerer significantly improves performance, regardless of whether explicit logic reasoning is employed.
Answering Logic RefCOCO RefCOCO+
67.2 70.3
80.8 76.9
70.8 73.1
83.3 78.4
Table 10: Contribution of components of NAVER. Results are based on the RefCOCO (IoU) and RefCOCO+ (IoU) test A set.
@@ 7 Ablation Studies for VLMs
We evaluate the impact of four different Vision-Language Models (VLMs) [24, 53, 28, 9] on the performance of NAVER. The results are shown in Table 11. All tested models achieve comparable results on both RefCOCO and RefCOCO+. Among the four VLMs, InternVL2-8B achieves slightly better performance with IoU scores of 83.3 on RefCOCO and 78.4 on RefCOCO+. However, the differences between the VLMs are marginal, indicating that the architecture of NAVER is insensitive to the choice of VLM. This shows NAVER’s ability to decompose complex tasks into simpler subtasks, reducing the reliance on the capabilities of the VLMs. Given the similar performance across all tested VLMs, we select InternVL2-8B for its availability and accessibility, ensuring simpler implementation without compromising performance. This demonstrates the flexibility of NAVER in adopting different VLMs while delivering state-of-the-art performance.
@@
BLIP2-XXL [24] 83.0 78.3
xGen-MM [53] 82.3 77.5
LLaVA1.5-7B [28] 82.9 78.4
InternVL2-8B [9] 83.3 78.4
Table 11: Comparison of VLM selection for NAVER. All results are evaluated on the RefCOCO and RefCOCO+ test A set.
@@ 9 Query Length Analysis
We analyze the impact of text query length on performance, comparing NAVER with baselines across different query lengths. The results, shown in Figure 4, indicate that NAVER consistently outperforms all baselines regardless of query length. While most methods experience a performance decline as query length increases, the extent of decrease varies. Notably, older models like GLIP [25] show a significant performance drop for longer queries, suggesting difficulties in handling complex text inputs. In contrast, NAVER maintains stable performance, showing its robustness in processing longer and more complex queries.
Figure 4: IoU Performance of different text query length in NAVER. The x-axis represents the length of text query, while the y-axis shows the IoU score of NAVER and baselines within each group of query length.

### v2

@@ 5 Conclusion
This supplementary material provides additional details about the proposed method NAVER. In particular, it provides details about the additional quantitative comparison result with both accuracy and IoU in section 6, the ablation studies for VLM in section 7, the analysis of self-correction mechanisms in section 8, the performance analysis for query length in section 9, and the prompts used for the LLMs and VLMs in section 10.
RefCOCO RefCOCO+ RefCOCOg Ref-Adv
Method Acc. IoU Acc. IoU Acc. IoU Acc. IoU
E2E GLIP-L [26] 55.0 54.1 51.1 51.3 54.6 54.8 55.7 55.2
KOSMOS-2 [38] 57.4 - 50.7 - 61.7 - - -
YOLO-World-X [10] 12.1 12.7 12.1 12.7 32.9 33.8 32.2 34.2
YOLO-Wrold-V2-X [10] 19.8 20.0 16.8 17.2 36.5 37.3 33.1 34.8
GroundingDINO-T [31] 61.6 60.2 59.7 58.9 60.6 59.7 60.5 59.8
GroundingDINO-B [31] 90.8 85.2 84.6 77.5 80.3 69.4 78.0 73.1
SimVG [12] 94.9 86.9 91.0 83.9 88.9 81.3 74.4 70.7
Florence2-B [52] 94.5 89.5 91.2 86.5 88.3 85.0 72.2 71.9
Florence2-L [52] 95.1 90.6 92.5 88.2 90.9 87.6 71.8 71.8
Compositional Code-bison [43] 44.4 - 38.2 - - - - -
ViperGPT† [44] 62.6 59.4 62.3 58.7 67.2 63.6 60.7 58.6
HYDRA† [21] 60.9 58.2 56.5 54.9 62.9 60.8 54.4 53.5
NAVER† 70.1 67.9 64.1 63.8 69.5 59.2 65.1 63.4
ViperGPT‡ 67.1 64.0 73.2 68.8 65.6 63.5 60.1 59.5
HYDRA‡ 63.5 61.3 62.9 60.9 59.8 58.4 53.2 53.5
NAVER‡ 91.7 83.3 82.4 78.4 75.9 72.5 87.3 74.2
ViperGPT* 68.6 66.5 73.8 70.4 68.7 66.8 58.2 58.3
HYDRA* 65.7 64.6 66.2 65.3 59.9 60.5 48.3 51.8
NAVER* 96.2 91.7 92.8 88.4 91.6 88.1 75.4 75.3
Table 10: Accuracy and IoU performance on the referring expression detection task. Results are shown on the RefCOCO, RefCOCO+, RefCOCOg [20], and Ref-Adv [1] datasets. In compositional methods, they are grouped with the same VLMs for fair comparison. The methods with same symbols (†, ‡, *) use the same VFMs. The VFMs used are: (†) uses GLIP-L and BLIP; (‡) uses GroundingDINO-B and InternVL2; (*) uses Florence2-L and InternVL2, respectively. All groups use GPT-4o Mini.
6 Additional Quantitative Comparison
In this section, we report both the acurracy and Intersection over Union (IoU) performance for the referring expression detection task. Table 10 presents a quantitative comparison on the RefCOCO, RefCOCO+, RefCOCOg [20], and Ref-Adv [1] datasets. Additionally, we include results for the compositional methods using GroundingDINO as the VFM, marked with ‡. We observed that for each configuration, our proposed method NAVER outperforms the other compositional methods using the same foundation models and also outperform the grounding methods themselves. These results further validate that our approach consistently achieves superior IoU performance over existing baselines.
@@ 7 Ablation Studies for VLMs
We evaluate the impact of four different Vision-Language Models (VLMs) [25, 53, 29, 9] on the performance of NAVER. The results are shown in Table 11. All tested models achieve comparable results on both RefCOCO and RefCOCO+. Among the four VLMs, InternVL2-8B achieves slightly better performance with accuracy scores of 96.2 on RefCOCO and 92.8 on RefCOCO+. However, the differences between the VLMs are marginal, indicating that the architecture of NAVER is insensitive to the choice of VLM. This shows NAVER’s ability to decompose complex tasks into simpler subtasks, reducing the reliance on the capabilities of the VLMs. Given the similar performance across all tested VLMs, we select InternVL2-8B for its availability and accessibility, ensuring simpler implementation without compromising performance. This demonstrates the flexibility of NAVER in adopting different VLMs while delivering state-of-the-art performance.
Figure 4: Accuracy performance of different text query length in NAVER. The x-axis represents the length of text query, while the y-axis shows the accuracy of NAVER and baselines within each group of query length.
@@
BLIP2-XXL [25] 95.4 91.7
xGen-MM [53] 94.5 87.9
LLaVA1.5-7B [29] 95.5 91.9
InternVL2-8B [9] 96.2 92.8
Table 11: Comparison of VLM selection for NAVER. All results (accuracy) are evaluated on the RefCOCO and RefCOCO+ testA datasets [20].
@@ 9 Query Length Analysis
We analyze the impact of text query length on performance, comparing NAVER with baselines across different query lengths. The results, shown in Figure 4, indicate that NAVER consistently reaches SoTA performance comparing to all baselines regardless of query length. While most baseline methods experience a performance decline as query length increases, the extent of decrease varies. Notably, older models like GLIP [26] show a significant performance drop for longer queries, suggesting difficulties in handling complex text inputs. In contrast, NAVER maintains stable performance, showing its robustness in processing longer and more complex queries.